The following fractions represent just three different numbers. Separate them into three groups of equivalent fractions, by changing each one to its simplest form.
8/50

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

8 / 50 = (4 × 2) / (25 × 2)
= 4 / 25

Thanks